Circular No. 8671

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA OPHIUCHI 2006 G. Pojmanski, Warsaw University Astronomical Observatory, reports

the discovery on ASAS-3 200-mm-f.l.-lens telephoto images (cf. IAUC 8617) of an apparent nova located at α = 17^{h}33^{m}13^{s}, δ = −24^{o}21′.1 (equinox 2000.0). Available ASAS V magnitudes: 2005 Oct. 26.013 UT, [14; 2006 Feb. 9.390, 12.08; 10.327, 12.04; 10.329, 11.94; 10.342, 12.07; 11.334, 11.48; 12.329, 11.07; 13.336, 11.47. The object is also clearly visible in the ASAS-3 I-band images taken on Feb. 8.379. H. Yamaoka, Kyushu University, provides his average of measurements of the position of the new nova from I_{c}-band CCD images taken by S. Kiyota on Feb. 12.83 and by H. Maehara on Feb. 10.83: α = 17^{h}33^{m}13^{s}.06, δ = −24^{o}21′07′′.1 (rms ∼ 0′′.2; UCAC2 reference stars); the nearest star in the USNO-B1.0 catalog is ∼ 3′′ away from this position, and a red Digitized Sky Survey image from 1991 Aug. 11 does not show any object to limiting mag ∼ 18 (affected by nearby stars). Yamaoka adds that low-resolution spectra of this new object — taken around Feb. 12.9 by K. Ayani with the Bisei Astronomical Observatory (BAO) 1.01-m telescope, by H. Naito and S. Narusawa (Nishi-Harima Astronomical Observatory) with the 2.0-m NAYUTA telescope, and by M. Fujii (Ibara, Okayama, Japan) with a 0.28-m telescope — show prominent Balmer lines with P-Cyg profiles superimposed on a very red continuum, which suggests that the object is a classical nova near maximum. Deep interstellar Na D absorption is also seen. From the BAO spectrum, the FWHM of Hα emission is 560 km/s, and an absorption minimum of that is blueshifted by 530 km/s from the emission peek. O I 777.3-nm and Fe II emissions are also detected with P-Cyg profiles.

RS OPHIUCHI This recurrent nova is in outburst for the first time since 1985 (cf. IAUC

4030, 4031), as indicated by the following visual magnitude estimates: Feb. 12.829 UT, 4.5 (H. Narumi, Ehime, Kita-gun, Japan; communicated by K. Hirosawa, Aichi, Inazawa-shi, Japan); 12.847, 4.6 (K. Kanai, Gunma, Isezaki-shi, Japan; via Hirosawa); 12.862, 4.4 (Kanai); 13.163, 4.8 (W. Renz, Karlsruhe, Germany); 13.27, 4.9 (A. Pereira, Cabo da Roca, Portugal). S. Nakano (Sumoto, Japan) forwards unfiltered CCD magnitudes of RS Oph from Y. Nakamura (Kameyama, Mie-ken, Japan; survey frames with a 105-mm f/2.8 lens): Feb. 9.855, 10.5; 12.8426, 4.5. ASAS V magnitudes from Pojmanski: 2005 Oct. 26.998, 10.72; 2006 Feb. 13.341, 5.4 (saturated); 13.344, 5.3.

2006 February 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green